OpenStudy (anonymous):

use factoring to solve this equation x^2+x=56

OpenStudy (anonymous):

You can rewrite it as\[x ^{2}+x-56=0\]Then,\[x ^{2}+8x-7x-56=0\]\[x(x+8)-7(x+8)=0\]\[(x+8)(x-7)=0\]Thus, x=-8,7